-5/7k= 25
what is k?

Answer by checkley79(3341)

-5/7k= 25
IF YOU MEAN -5K/7=25 THEN:
-5K=25*7
-5K=175
K=175/-5
K=-35 ANS.
IF YOU MEAN -5/(7K)=25 THEN:
25*7K=-5
175K=-5
K=-5/175
K=-1/35 ANS.